# Real-Time Incremental Explanations for Object Detectors

### v1

@@ 3.3 Sufficient explanations
Thus far, our focus has been primarily on saliency maps. However, in addition to this, it is useful to provide sufficient explanations. A sufficient explanation, informally, is a subset of pixels capable of producing the desired classification. This is a useful sanity check on the saliency maps: the “hotspots” on the saliency map should be enough to trigger the appropriate classification. This necessitates the introduction of an “Explanation” block for both the initial frame and subsequent frames, as illustrated in Figure 2. This implementation ensures that the explanations provided for subsequent frames are correct, meaning they effectively output the intended classification with a bounding box containing an area with the correct classification.
Extraction of sufficient explanations from saliency maps can be slow. Therefore we adopt the approach used in Algorithm 2. For the initial frame 00, the landscape is divided into $l$ equal sections, demarcated by grey dotted lines in Figure 3(a). To efficiently locate an explanation, we employ binary search within these predefined divisions. In Figure 3(a), the division containing the explanation threshold is highlighted in red. We assume monotonicity in the model output as pixels are added: once a sufficient set of pixels has been located, the classification no longer changes.
Explanation discovery for subsequent frames is depicted in Figure 3(b). Assuming consistency in the saliency map, it is reasonable to expect that the explanation threshold in subsequent frames would be close to that of the previous frames. Therefore we search at around the previous threshold, with the number of divisions adjusted to $l_{n}$. Smaller values of $l_{n}$ improve performance but may result in explanations that contain more redundancy.
@@
Figure 3: Visualization of the explanation extraction procedure.
Algorithm 2 Explanation Algorithm
Input: Saliency map $s$, object detector $g$, image $I$, time step $t$, margin $\delta$, number of divisions $l$ (for $t=0$) and $l_{n}$ (for $t>0$)
Output: Explanation $xp$
1:  if $t==0$ then
2:     $\text{min}\leftarrow\min(s)$
3:     $\text{max}\leftarrow\max(s)$
4:     $ss\leftarrow\text{Linspace}(\text{min},\text{max},l)$
5:  else
6:     $\text{max}\leftarrow\text{threshold}_{t-1}\cdot(1+\delta)$
7:     $\text{min}\leftarrow\text{threshold}_{t-1}\cdot(1-\delta)$
8:     $ss\leftarrow\text{Linspace}(\text{min},\text{max},\l_{n})$
9:  end if
10:  $xp,\text{threshold}_{t-1}\leftarrow\text{BinarySearch}(s,g,I,ss)$
11:  return $xp$
Algorithm 2 provides a pseudocode of this algorithm. For the first frame, the search space $ss$ spans from the minimum to maximum values of the saliency map $s$, divided into $\l$ equally spaced levels. For subsequent frames, an additional parameter $\delta$ defines the margin around the previous $\text{threshold}_{t-1}$, adjusting the search space boundaries accordingly and setting the number of divisions to ł$\l_{n}$. A binary search is performed within the defined search space for each case to determine the explanation.
Model Explainer Ins. ($\uparrow$) Del. ($\downarrow$) EPG ($\uparrow$) EP ($\downarrow$) Time ($\downarrow$)
yolo v10 d-rise $0.6748$ $0.0226$ $0.2744$ $0.2868$ $30.0458$
incx $0.7231$ $0.0211$ $0.2565$ $0.2627$ $0.7692$
faster r-cnn d-rise $0.8975$ $0.0164$ $0.2449$ $0.0465$ $40.2852$
incx $0.8891$ $0.0211$ $0.2693$ $0.0568$ $0.3645$
rt-detr d-rise $0.8993$ $0.0327$ $0.2448$ $0.0547$ $57.7377$
incx $0.8854$ $0.0374$ $0.2501$ $0.0514$ $0.6668$
Table 1: Comparison of different models using d-rise and incx with d-rise for the first explanation with their respective metrics. The arrows indicate whether higher $\uparrow$ or lower $\downarrow$ values of each metric are better.
@@ 4 Experimental Results
The COCO (Lin et al. 2014) and VOC (Everingham et al. 2010) datasets are typically used for evaluation of explainability in object detection (Kuroki and Yamasaki 2024). However, these datasets consist of independent images that lack temporal continuity, making them unsuitable for our focus on real-time video explanation. LaSOT (Fan et al. 2019), however, includes 85 object classes and over 3.87 million frames in total, which is more appropriate for our experiments. Therefore, we curated a subset comprising two object classes: car and cat. From these classes, we selected the first $300$ frames (equivalent to 10 seconds at $30$ FPS) from a total of $10$ videos, resulting in $3000$ frames for our experiments.
@@
We executed incx, using d-rise to compute the saliency maps, to be able to perform a comparison between incx and d-rise. incx makes no assumptions about how the initial saliency map is produced: we use d-rise as it works naturally with the different object detectors that we examine. We applied incx to all the videos, tracking the object with the highest confidence from the first frame. We then computed the saliency maps using d-rise for all frames where the object was successfully tracked by incx. All experiments were conducted using the NVIDIA A100 Tensor Core GPU. We executed d-rise with $1000$ mutants and a mask resolution of (4,4). The following versions of the models were used: yolov10-n, rt-detr-l, and the PyTorch implementation of faster r-cnn.
Figure 4: Typical normalized insertion and deletion curves. Model confidence increases quickly as pixels are added, but then largely flattens. The deletion curve’s rapid descent to 00 is due to the disappearance of the object from the perspective of the object detector: the bounding box vanishes.
@@
For a one-to-one comparison between each saliency map generated by d-rise and incx, we computed a set of metrics commonly used in the field of Explainable AI. Below, we introduce these metrics.
Normalized Insertion/Deletion curves
@@
The insertion process starts with a blank image, adding pixels based on the saliency map, and computes the area under the curve (AUC) after measuring detector confidence. Conversely, the deletion process starts with the full image and gradually removes pixels according to the saliency map. As we are dealing with object detectors, we adjust this metric by multiplying the score by the Intersection over Union (IoU) of the original bounding box. To account for low confidence producing low insertion and deletion values, we normalize the curve by dividing all values by the original image’s score. Figure 4 shows typical curves for both metrics.
Energy-based Pointing Game (EPG)
@@
The Energy-based Pointing Game (EPG) metric quantifies the proportion of the saliency map’s total energy that is contained within the bounding box (Wang et al. 2019). This metric is essential because an accurate saliency map should predominantly cover the bounding box to correctly represent the object of interest. If the saliency map is focused outside of the bounding box, it may indicate the detection of an object that does not correspond to the one within the current bounding box.
Explanation Proportion (EP)
@@
The Explanation Proportion metric measures explanation quality by computing the ratio of the explanation to the whole image, hence favoring compact explanations that minimize unnecessary information. We define EP as follows: for a given boolean-valued mask $m$ of an image, where $\top$ denotes the pixels included in the explanation, the Explanation Proportion is computed as:
@@
As summarized in Table 1, the insertion values for incx are consistently within 8% of those observed for d-rise, with all other metrics displaying similar trends. Moreover, the computation time for incx is two orders of magnitude lower than that of d-rise ($111$ times faster for faster r-cnn), highlighting the suitability of incx for real-time scenarios.
Saliency maps Explanations
Model CC SSIM JI DC
yolo v10 $0.7909$ $0.6518$ $0.4713$ $0.5931$
faster r-cnn $0.8517$ $0.6746$ $0.4481$ $0.5878$
rt-detr $0.8592$ $0.7110$ $0.4546$ $0.5910$
Table 2: Similarity of saliency maps and explanations for different models using d-rise and incx
Given that incx generates an approximation of the true saliency map, it is crucial to evaluate the similarity between the baseline saliency map and the approximation provided by incx. To achieve this, we need to employ metrics that effectively measure the similarity between these two saliency maps. Mean Squared Error (MSE) is not ideal for this purpose, as it may not provide a clear indication of perceptual similarity between the maps. Instead, we utilize alternative metrics designed to assess the correspondence between the baseline and incx saliency maps. The selected metrics for this comparison are the Pearson Correlation Coefficient (CC), Structural Similarity Index (SSIM) (Wang et al. 2004), Dice Coefficient (DC) (Dice 1945), and Jaccard Index (IoU). The results presented in Table 2 show a strong similarity between the approximations implemented by incx and the original saliency maps created by d-rise.
@@
incx is based on the assumption that the object it is applied to is rigid. Moreover, the algorithm may not handle rotations (pitch, roll, and yaw) effectively, especially for non-symmetrical objects, given that it primarily addresses scaling and translation. Therefore, the method may not be theoretically advantageous for objects undergoing other types of transformations. Nonetheless, as shown in Section 4, these theoretical limitations have a limited impact in practice. Furthermore, to ensure the accurate computation of explanations in real-time, we assume that explanations for consecutive frames are relatively close on the saliency map, as shown in Figure 3. However, this approach may result in non-minimal explanations.
@@ 5 Conclusions
Motivated by the need for real-time explanations for object detectors, we have introduced incx, an algorithm and tool for approximating saliency maps via linear transformations. Our results demonstrate that incx operates in real time, with negligible overhead over the object detector and is two orders of magnitude faster than generating a new saliency map. Moreover, incx produces saliency maps closely aligned with freshly computed maps, achieving an average correlation coefficient of approximately $0.8$ across all models used, with minimal computational overhead. The quality metrics for assessing explainable AI algorithms show that incx is comparable to d-rise, with the insertion score within $8\%$ of d-rise’s ones.
incx handles linear transformations of objects; future work will address the current limitations related to rotation and object deformation. We anticipate future applications of this algorithm in various industries, particularly in self-driving cars and video advertisement placement.

### v2

@@ 3.3 Sufficient explanations
Thus far, we have discussed saliency maps. It is also useful to provide sufficient explanations. We use the following definition, which is a standard definition of causal explanations for image classifiers [8].
Definition 3 ([8]).
An explanation is a minimal subset of pixels of a given input image that is sufficient for the model to correctly detect the object, with all other pixels of the image set to a baseline value (occluded).
Computing minimal and sufficient explanations is intractable [8], so we relax the minimality requirement. A sufficient explanation is a subset of pixels of the input image that is sufficient for the AI model to reproduce the original classification of the entire image. This is a useful sanity check of the saliency maps: the “hot spots” on the saliency map should be sufficient to trigger the original object detection. Sufficient explanations are computed by the explanation procedure for both the initial and the subsequent frames, as shown in Figure 2. We note that, while the saliency maps for subsequent frames are approximate, the extracted explanations are still correct, that is, sufficient for the original classification. They can, however, be larger than those computed from baseline saliency maps.
A brute-force extraction of sufficient explanations from saliency maps is inefficient, as it requires checking all subsets of pixels of the image. Our Explanation procedure, illustrated in Figure 3(b) (see also the pseudo-code in the supplementary material) uses a binary search on the saliency map of the current frame, as described below.
For the initial frame, the saliency map is divided into $l$ equal horizontal sections, as illustrated in Figure 3(a) (lateral view). The goal is to determine the smallest region that is sufficient for the model to make the same prediction. Therefore, we progressively occlude the pixels under a given dotted line. We then query the model with the partially occluded input to check whether the object is still being detected.
A naive approach would require querying the model at every possible dotted line, which is computationally expensive. Instead, we use binary search to efficiently locate a threshold (depicted by the red dotted line) that provides a sufficient, albeit not necessarily minimal, explanation. This approach works because the property of being sufficient is monotonic. In other words, starting from the top of the saliency map, as we reveal more of the image, the explanation becomes sufficient at some point and stays sufficient for all subsequent values.
Explanation discovery for subsequent frames is depicted in Figure 3(b). Assuming temporal consistency in the saliency map, it is reasonable to expect that the explanation threshold in subsequent frames will be close to that of the previous frames. Therefore, we start our binary search around the previous threshold, with the number of divisions adjusted to a parameter $l_{n}$. Smaller values of $l_{n}$ improve performance but may result in explanations that contain more redundancy.
@@
Figure 3: Visualization of the Explanation procedure.
Dataset Model Explainer Ins. ($\uparrow$) Del. ($\downarrow$) EPG ($\uparrow$) EP ($\downarrow$) Time(s) ($\downarrow$)
BDD100K Faster R-CNN D-RISE $0.9459$ $0.0233$ $0.1974$ $0.0125$ $170.55$
IncX $0.9448$ $0.0246$ $0.2009$ $0.0151$ $1.06$
RT-DETR D-RISE $0.9321$ $0.0510$ $0.2542$ $0.0387$ $130.30$
IncX $0.9312$ $0.0478$ $0.2658$ $0.0304$ $2.54$
YOLO D-RISE $0.7879$ $0.0326$ $0.2771$ $0.0816$ $44.83$
IncX $0.8109$ $0.0354$ $0.3159$ $0.0700$ $1.22$
KITTI Faster R-CNN D-RISE $0.9705$ $0.0212$ $0.0968$ $0.0070$ $129.06$
IncX $0.9646$ $0.0214$ $0.1384$ $0.0074$ $2.15$
RT-DETR D-RISE $0.9741$ $0.0139$ $0.0818$ $0.0114$ $114.09$
IncX $0.9760$ $0.0157$ $0.1071$ $0.0131$ $4.23$
YOLO D-RISE $0.9031$ $0.0153$ $0.1197$ $0.0485$ $40.32$
IncX $0.8970$ $0.0158$ $0.1330$ $0.0521$ $2.60$
NuScenes Faster R-CNN D-RISE $0.9827$ $0.0035$ $0.0322$ $0.0089$ $203.56$
IncX $0.9771$ $0.0034$ $0.0366$ $0.0118$ $23.25$
RT-DETR D-RISE $0.9119$ $0.0304$ $0.1199$ $0.0293$ $225.65$
IncX $0.9011$ $0.1182$ $0.1201$ $0.0272$ $38.84$
YOLO D-RISE $0.7639$ $0.0117$ $0.1212$ $0.0813$ $95.45$
IncX $0.7777$ $0.0132$ $0.1202$ $0.1178$ $14.37$
VIPER Faster R-CNN D-RISE $0.9417$ $0.0452$ $0.1705$ $0.0177$ $277.44$
IncX $0.9368$ $0.0641$ $0.1502$ $0.0094$ $22.91$
RT-DETR D-RISE $0.9465$ $0.1420$ $0.3373$ $0.0617$ $245.32$
IncX $0.9388$ $0.1304$ $0.4059$ $0.0777$ $27.91$
YOLO D-RISE $0.7128$ $0.0555$ $0.3284$ $0.1280$ $114.30$
IncX $0.7482$ $0.0599$ $0.4406$ $0.1154$ $9.80$
Table 1: Comparison of different models across multiple datasets using d-rise and IncX as explainers. The arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values of each metric are better.
@@ 4 Experimental Results
Dataset Model CC SSIM
BDD100K Faster R-CNN $0.9395$ $0.8599$
RT-DETR $0.9648$ $0.8383$
YOLO $0.9307$ $0.7447$
KITTI Faster R-CNN $0.9071$ $0.5832$
RT-DETR $0.9118$ $0.5625$
YOLO $0.9308$ $0.7471$
NUSCENES Faster R-CNN $0.9433$ $0.8284$
RT-DETR $0.9028$ $0.8023$
YOLO $0.9221$ $0.8494$
VIPER Faster R-CNN $0.9147$ $0.8308$
RT-DETR $0.8661$ $0.7124$
YOLO $0.8203$ $0.6464$
Table 2: Correlation Coefficient (CC) and Structural Similarity Index Measure (SSIM) scores used for comparing saliency maps across different models for each dataset.
Dataset Model JI DC
BDD100K Faster R-CNN $0.3308$ $0.4371$
RT-DETR $0.5582$ $0.6611$
YOLO $0.6660$ $0.7853$
KITTI Faster R-CNN $0.4323$ $0.5535$
RT-DETR $0.5655$ $0.6996$
YOLO $0.6861$ $0.7993$
NUSCENES Faster R-CNN $0.5603$ $0.6774$
RT-DETR $0.5387$ $0.6325$
YOLO $0.7144$ $0.7967$
VIPER Faster R-CNN $0.4220$ $0.5055$
RT-DETR $0.4252$ $0.5466$
YOLO $0.6147$ $0.7380$
Table 3: Jaccard Index (JI) and Dice Coefficient (DC) scores used for comparing explanations across different models for each dataset.
The COCO [29] and VOC [14] datasets are typically used for evaluation of explainability in object detection [26]. However, these datasets consist of independent images that lack temporal continuity, making them unsuitable for our focus on real-time video explanation for autonomous vehicles. To address this limitation, we selected datasets specifically designed for autonomous driving, namely BDD100K [47], KITTI [16], NuScenes [6] and VIPER [40], each of which contain hundreds of thousands of individual frames distributed across hundreds of videos. Given the computational expense of explanation methods such as d-rise, processing these datasets in full would be impractical. To ensure feasible benchmarking, we randomly selected a representative subset from each dataset. From KITTI, BDD100K, and VIPER, we chose five videos each and extracted 100 frames per video. For KITTI, this corresponds to 10 seconds per video at 10 FPS. For NuScenes, we used the entire mini dataset, consisting of 10 videos of 404 frames.
@@
We evaluated IncX on the following object detectors: yolov10-n, rt-detr-l, and the PyTorch implementation of faster r-cnn. d-rise was used to compute the initial saliency maps. IncX makes no assumptions about how the initial saliency map is produced: d-rise is used as it works naturally with the different object detectors that we examine and is considered the state-of-the-art explainability tool for object detectors. For comparison we also computed the saliency maps independently using d-rise for all frames. We executed d-rise with $1000$ mutants and a mask resolution of $(4,4)$. All experiments were conducted using the NVIDIA A100 Tensor Core GPU.
@@
For a one-to-one comparison between each saliency map generated by d-rise and IncX, we computed a set of metrics commonly used in the field of Explainable AI. We describe these metrics below.
@@
The Normalized Insertion/Deletion curves metric starts with a blank image, adding pixels based on the saliency map, and computes the area under the curve (AUC) after measuring detector confidence. Conversely, the deletion process starts with the full image and gradually removes pixels according to the saliency map. For object detectors, we adjust this metric by multiplying the score by the Intersection over Union (IoU) of the original bounding box. This adjustment accounts for the fact that traditional insertion and deletion curves assess whether an object is detected, but do not consider its precise location, an essential factor for object detection tasks. To account for low confidence producing low insertion and deletion values, we normalize the curve by dividing all values by the original image’s softmax classification.
@@
The Energy-based Pointing Game (EPG) metric quantifies the proportion of the saliency map’s total energy that is contained within the bounding box [44]. This metric is essential because an accurate saliency map should predominantly cover the bounding box to correctly represent the object of interest. If the saliency map is focused outside of the bounding box, it may indicate the detection of an object that does not correspond to the one within the current bounding box.
@@
The Explanation Proportion (EP) metric measures explanation quality by computing the ratio of the explanation to the whole image, hence favoring compact explanations that minimize unnecessary information. We define EP as follows: for a given boolean-valued mask $m$ of an image, where $\top$ denotes the pixels included in the explanation, the EP is computed as
@@
As summarized in Table 1, the insertion values for IncX are consistently within 5% of those observed for d-rise, with all other metrics displaying similar trends. Moreover, the running time of IncX is two orders of magnitude less than that of d-rise in most cases, showing that IncX is suitable for real-time scenarios.
The quality of an approximate saliency map computed by IncX is evaluated by measuring its deviation it from the baseline saliency map, computed by d-rise directly from the image. We compute this deviation by using standard metrics: Pearson Correlation Coefficient (CC), Structural Similarity Index (SSIM) [45] (Table 2), Dice Coefficient (DC) [11], and Jaccard Index (IoU) (Table 3). The results in both tables indicate a strong similarity between the saliency maps generated by IncX and the baseline maps produced by d-rise.
@@
IncX assumes that the object it is applied to is rigid and does not rotate. Our experimental results show that this assumption is reasonable in practice for autonomous driving. Another implicit assumption is that consecutive frames produce similar saliency maps, which is supported by the laws of physics: objects tend not to disappear or (re-)materialize.
@@ 5 Conclusions
In this paper we have presented the algorithm IncX for real-time explanations of object detectors. The algorithm incurs a negligible overhead on the object detector and is two orders of magnitude faster than the state-of-the-art. Our method does not lead to inferior quality saliency maps, despite its assumptions. The standard quality metrics for assessing XAI algorithms show that IncX is comparable to d-rise, the state-of-the-art in black-box XAI for object detectors. Evaluation across multiple standard datasets supports our claim that IncX is suitable for autonomous driving applications.